A 3 m^3 container is filled with 6 kg of steam at 0.5 MPa. Describe its phase and find the temperature, specific and total volume, specific and total internal energy, specific and total enthalpy, and specific and total entropy.
Answer #1

Solution: Given Total volume, that - V-3m m= 6kg. P=0.5 mla=star. specific volume, & av = 2 20.50 m/reg. from Steam Table (saSpecific enthalpy ext of steam at 276.55°c. ha 2964+ (30648 –29411) weds = 296.1 + ( 3064.8-2961.11 1 x26.55 50 h = 3016116 kTotal internal energy of the steam Uzmiu = 6*2766-16 kJ - 16.597 MJ.